# Chaos and spatial prethermalization in driven-dissipative bosonic chains

### v2

@@ II.2 Regular and chaotic regimes in the NESS
which quantifies the relative distance to Poissonian statistics  [^1] Compared to the conventional indicator $g^{(2)}$, $\delta n^{\rm ss}_{\ell}$ defined in Eq. (6) is less prone to numerical errors caused by a finite sampling of Wigner trajectories.. $\langle...\rangle_{\rm ss}$ indicates the average once the NESS is reached. $\delta n^{\rm ss}_{\ell}=0$ corresponds to Poissonian statistics typical of coherent states, $\delta n^{\rm ss}_{\ell}>0$ corresponds to super-Poissonian statistics, and sub-Poissonian statistics with $\delta n^{\rm ss}_{\ell}<0$ are known to be incompatible with a classical description of the state [61].
@@
Focusing on the last site, $\ell=L$, the results are collected in Fig. 3. Qualitatively similar results are obtained throughout the chain, and we refer the reader to the Supplementary Information for results at $\ell=1$ and $\ell=L/2$. We also provide representative instances of local Wigner functions $W_{L}^{\rm ss}(\alpha,\alpha^{*})$. We identify three distinct regimes:
Figure 4: Spatial relaxation of amplitude and phase degrees of freedom in the chaotic regime. Spatial profiles of equal-time statistics of the complex field, decomposed as $\alpha_{\ell}=:r_{\ell}\,{\rm e}^{{\rm i}\phi_{\ell}}$ with $\ell=1,...L$, once the chaotic NESS is reached.
(a) Average amplitude $\langle r_{\ell}\rangle_{\rm ss}$, showing a large and increasing relaxation scale for increasing chain lengths $L=$ 80, 200, 300, 400.
(b) Average phase $\langle\phi_{\ell}\rangle_{\rm ss}$, rapidly converging to zero. (c) First-order coherence function $|g_{k,\ell}^{(1)}|$ defined in Eq. (7) for $k=L$ (blue) and $k=L/2$ (light-blue), showing sharp exponential decorrelation of the phases away from $\ell=k$. The drive strength is set to $F=7.5$ and the other parameters are set as in Fig. 2.
@@
    Regular quasilinear regime; at weak drive, only few photons populate the last site. The photon statistics are Poissonian ($\delta n^{\rm ss}_{L}\approx 0$) and the small saturation value of the phase OTOC, $D^{\rm ss}_{1,L}(\tau\to\infty)\approx 10^{-2}-10^{-1}$, indicates regular dynamics. The Wigner function exhibits the Gaussian shape characteristic of a coherent state. In this regime, single-particle excitations are dilute, rendering nonlinearities negligible and preventing the onset of chaos.
@@
    Chaotic regime; at stronger drives, the photon number markedly increases with respect to the quasilinear regime and nonlinearities are now relevant. The large value of $\delta n^{\rm ss}_{L}$ indicates strong super-Poissonian fluctuations. Concurrently, the phase OTOC saturates to $D^{\rm ss}_{1,L}(\tau\to\infty)\approx 1$, signaling significant dephasing between the fields in the two replicas. This is also reflected in the local Wigner function where the ergodicity of Wigner trajectories translates into $W_{L}^{\rm ss}(\alpha,\alpha^{*})$ spreading across a wide portion of phase space. These observations, along with the exponential growth of $D_{1,\,\ell}^{\rm ss}(\tau)$ shown in Fig. 2 (a) for a representative value of $F$ indicated by a pentagon marker, confirm the chaotic nature of the quantum dynamics.
@@
    Regular coherent regime; at even stronger drives, the photon number statistics become sub-Poissonian ($\delta n_{L}<0$). This regime is incompressible in the sense that the photon number marginally increases with $F$ [62]. The Wigner function acquires a crescent-shaped distortion with respect to the coherent state [44] which is a typical signature of many-body phases driven by the competition of Kerr-like nonlinearities and detuning, such as in optical bistability [63]. The small saturation value $D_{1,\,L}^{\rm ss}(\tau\to\infty)\approx 10^{-2}-10^{-1}$ indicates persistent phase correlations between the two replicas. These observations, along with the sub-exponential growth of $D_{1,\,\ell}^{\rm ss}(\tau)$ shown in Fig. 2 (b) for a representative value of $F$ indicated by a square marker, are indicative of regular dynamics.
Notably, the crossover from the regular quasilinear regime (I) to the chaotic regime (II) is smooth in both $n^{\rm ss}_{L}$ and $\delta n^{\rm ss}_{L}$ while the transition between the chaotic regime (II) and the regular coherent regime (III) is characterized by abrupt variations. In particular, the drive strength $F$ at which the photon number leaps coincides with that of the maximum of $\delta n^{\rm ss}_{L}$.
II.3 Two-stage thermalization in space
In this section, we focus on the chaotic regime by fixing the drive strength to $F=7.5$. We decompose the complex field in terms of amplitude and phase degrees of freedom, $\alpha_{\ell}=r_{\ell}\,{\rm e}^{{\rm i}\phi_{\ell}}$ where $r_{\ell}\geq 0$, and we separately analyze their spatial relaxation along the chain, from $\ell=1$ to $\ell=L$.
Figure 4 shows the steady-state profile of the average amplitude $\langle r_{\ell}\rangle_{\rm ss}$ and average phase $\langle\phi_{\ell}\rangle_{\rm ss}$ across the chain for various chain lengths $L$. The amplitude of the field slowly decays along the chain. This relaxation is characterized by a length scale $\xi_{r}$ that clearly grows monotonically with the system size $L$. This large length scale can be explained by the local conservation of the photon number within the bulk of the chain, which hinders the rapid relaxation of the amplitude degree of freedom. Instead, this relaxation occurs through much slower hydrodynamic processes involving, notably, the driven-dissipative conditions at the two boundaries of the chain. In contrast, the phase degree of freedom averages to zero over a much shorter, microscopic, length scale $\xi_{\phi}$ spanning only a few sites.
@@ II.3 Two-stage thermalization in space
$g^{(1)}_{k,\ell}:=\frac{\langle\hat{a}_{k}^{\dagger}\hat{a}_{\ell}\rangle_{\rm ss}}{\sqrt{\langle\hat{a}_{k}^{\dagger}\hat{a}_{k}\rangle_{\rm ss}\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle_{\rm ss}}}\,.$ (7)
@@
The results are presented in Fig. 4 (c) as a function of $\ell$, for both $k=L/2$ and $k=L$. In both cases, the profile of the spatial correlations shows exponential decay, $|g^{(1)}_{k,\ell}|\sim{\rm e}^{-|k-\ell|/\xi_{\phi}}$, typical of a disordered phase with a correlation length $\xi_{\phi}$ on the order of a few sites, before saturating to values below $10^{-2}$.
@@
Altogether, these results suggest a two-stage thermalization process in space: the phase sector thermalizes to a disordered state over a short length scale, while the amplitude sector is contingent upon both boundary conditions and relaxes over the entire system. As we shall demonstrate below, this opens up a significant region of space, beyond the first few sites near the boundary drive, but before the complete relaxation to near vacuum states subject to thermal fluctuations, where the U(1) symmetry of the undriven Hamiltonian is restored and the highly-populated sites along the chain can be seen as in local equilibria characterized by a large and slowly-varying chemical potential.
@@
Figure 5: Prethermal phase in the chaotic regime. (a) Spatial profile of the quantity $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ across a chain of length $L=400$. $\beta^{\rm eff}_{\ell}$ and $\mu^{\rm eff}_{\ell}$ are the effective local inverse temperature and chemical potential, respectively. The results are obtained by mapping the local physics at site $\ell$ to the 2-photon impurity ansatz defined in Eqs. (8-9) and the quantity $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}$ is computed via Eq. (10). We also report the results obtained by mapping to the generalized Scully-Lamb ansatz and to an equilibrium Gibbs measure. Three distinct domains are discussed in the text: 1) nonsymmetric nonthermal domain where the ansatz fails to reproduce $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$ such as the one in panel (c) ; 2) prethermal domain with $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}>0$ ; 3) thermal domain with $\beta^{\rm eff}_{\ell}\mu^{\rm eff}_{\ell}\leq 0$. (b) Respective sizes of the three domains versus the overall chain length $L$. (c-e) Normalized local Wigner functions, $\tilde{W}_{\ell}^{\rm ss}(\alpha,\alpha^{*}):=W^{\rm ss}_{\ell}(\alpha,\alpha^{*})/\textrm{max}[W^{\rm ss}_{\ell}(\alpha,\alpha^{*})]$ for representative sites in the three domains of a chain of length $L=400$ and corresponding to the markers in panel (a). The drive strength is set to $F=7.5$ and the other parameters are set as in Fig. 3.
@@
We test this scenario by probing the local symmetries, the thermodynamics, and the temporal fluctuations across the chain. On general grounds, the local steady-state density matrix $\hat{\rho}^{\rm ss}_{\ell}:=\lim\limits_{t\to\infty}\textrm{Tr}_{k\neq\ell}\hat{\rho}(t)$ can be cast as the solution of an effective driven-dissipative impurity model constructed by singling out one site of the chain, and tracing out its neighbors such as to treat them as incoherent sources and drains. In perfect thermal equilibrium with U(1) symmetry, the statics of the impurity model are to be straightforwardly described by the Gibbs state $\hat{\rho}^{\rm eq}_{\ell}\propto\exp\small[-\beta(\hat{H}_{\rm imp}-\mu\,\hat{a}^{\dagger}a)]$ where $\beta$ and $\mu$ are the inverse temperature and chemical potential, respectively, and the impurity Hamiltonian
$\hat{H}_{\rm imp}=\omega_{0}\,\hat{a}^{\dagger}\hat{a}+\frac{1}{2}U\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}\,,$ (8)
corresponds to a single site of the chain Hamiltonian in Eq. (2) (see the Supplementary Information). In a generic NESS, however, the rigorous derivation of an effective driven-dissipative impurity model is a daunting task, and we simply rely on symmetries and heuristics to identify a minimal impurity model that successfully locally captures both the statics as well as temporal fluctuations. Assuming that the U(1) symmetry is indeed rapidly restored in the chain, this naturally brings us to consider the impurity ansatz defined by the local U(1)-symmetric Hamiltonian $\hat{H}_{\rm imp}$ in Eq. (8) and by the first non-trivial jump operators allowed under weak U(1) symmetry [64]
$\hat{L}_{\uparrow}\!=\!\sqrt{\gamma^{\uparrow}}\hat{a}^{\dagger}\!,\hat{L}_{\downarrow}\!=\!\sqrt{\gamma^{\downarrow}}\hat{a},\hat{L}_{\phi}\!=\!\sqrt{\gamma^{\phi}}\hat{a}^{\dagger}\hat{a},\hat{L}_{\rm s}\!=\!\sqrt{\gamma^{\rm s}}\hat{a}^{2}.$ (9)
The non-negatives parameters $\gamma^{\uparrow}$, $\gamma^{\downarrow}$, $\gamma^{\phi}$, and $\gamma^{\rm s}$ are effective rates of incoherent pumping, decay, dephasing, and 2-photon decay. The inclusion of the latter is essential to ensure the saturation of the photon number when $\gamma^{\uparrow}>\gamma^{\downarrow}$.
@@
We test the validity of this effective modeling, and the underlying restoration of the U(1) symmetry, by fitting the local steady-state Wigner function $W_{\ell}^{\rm ss}(\alpha,\alpha^{*})$ along the chain to the ones predicted by the impurity ansatz. The fitting parameters are $\gamma^{\uparrow}$, $\gamma^{\downarrow}$ and $\gamma^{\rm s}$ since the steady-state Wigner function of the impurity ansatz is independent of the Hamiltonian parameters and of the dephasing rate $\gamma^{\phi}$. The latter can be extracted by fitting two-time correlation functions. The overall fitting procedure is detailed in the Supplementary Information. We repeat this procedure at all sites from $\ell=1$ to $\ell=L$. Excellent matches are obtained everywhere along the chain except for the first few sites, confirming the scenario of the U(1) symmetry being rapidly restored away from the boundary drive.

### v3

@@ II.2 Dynamical regimes in the NESS
which quantifies the relative distance to Poissonian statistics  [^1] Compared to the conventional indicator $g^{(2)}$, $\delta n_{\ell}$ defined in Eq. (5) is less prone to numerical errors caused by a finite sampling of Wigner trajectories.. delimited-⟨$\langle$..$\rangle$ … ⟩ indicates the average over Wigner trajectories once the NESS is reached. $\delta n_{\ell}=0$ corresponds to Poissonian statistics typical of coherent states, $\delta n_{\ell}>0$ corresponds to super-Poissonian statistics, and sub-Poissonian statistics with $\delta n_{\ell}<0$ are incompatible with a classical description of the state [73]. The quantity $\delta n_{\ell}$ is directly related to the second-order coherence function $g^{(2)}_{\ell}$, a standard observable in quantum optics, via $\delta n_{\ell}=n_{\ell}^{2}(g^{(2)}_{\ell}-1)$.
In addition to the above static diagnostics, we probe the chaotic character of the dynamics, or lack thereof, by means of an out-of-time correlator (OTOC) in the NESS. OTOCs are commonly used as a diagnostic of the spatiotemporal spread of chaos in both quantum mechanical and classical systems. We focus on the OTOC between the number and the phase degrees of freedom of the resonators. In the Methods, we show that the semiclassical formulation of this phase OTOC reads [74, 75, 76, 61, 77, 78, 79]
$D_{k,\ell}(\tau):=1-\lim_{t\to\infty}\left\langle\cos\left[\varphi_{\ell}^{(a)}(t+\tau)-\varphi_{\ell}^{(b)}(t+\tau)\right]\right\rangle.$ (6)
Here, $\varphi_{\ell}(t):=\textrm{arg}[\alpha_{\ell}(t)]$ is the phase of the complex field in the $\ell$-th resonator along an individual Wigner trajectory. The superscripts $(a)$ and $(b)$ refer to two replicas of the system which are identical until time $t$, at which an infinitesimal perturbation is applied to the resonator at site $k$ in replica $b$ according to $\varphi_{k}^{(b)}(t)=\varphi_{k}^{(a)}(t)+\varepsilon$, with $\varepsilon\ll 1$. The subsequent evolution is computed using the same quantum noise realization for the two replicas. We stress that in Eq. (6), the replica (b) is the one perturbed at site $k$ and, therefore, $\varphi_{\ell}^{(b)}$ implicitly depends on the $k$ index.
Figure 2: Nonequilibrium steady-state phase diagram. Steady-state properties of the last resonator in the chain, $\ell=L$, as functions of the chain length $L$ and drive strength $F$. (a) Photon number $n_{L}$. (b) Photon-number fluctuation $\delta n_{L}$, defined in Eq. (5). (c) Saturation value of the steady-state phase OTOC, $D_{1,L}(\tau\to\infty)$, defined in Eq. (6). The three distinct regimes labeled I, II, and III are discussed in the text. Panels (d), (e), and (f) show cuts of panels (a), (b), and (c), respectively, at fixed chain length $L=10$. Results are computed by averaging over $N_{\rm traj}=10^{3}$ independent Wigner trajectories. Statistics are further improved by averaging over a time window $\Delta\tau$ after reaching the NESS: $\Delta\tau=25$ for $D_{1,L}(\tau\to\infty)$, and $\Delta\tau=2\times 10^{3}$ for $n_{L}$ and $\delta n_{L}$. Throughout the manuscript, the dissipation rate $\gamma$ sets the unit of energy, and the other parameters are fixed to $\Delta=2.5$, $J=2$, and $U=0.1$.
Both the initial growth and the late regimes of $D_{k,\ell}(\tau)$ shed light on the chaotic versus regular nature of the dynamics in the NESS. In chaotic dynamics, the typical distance between two trajectories with nearly identical initial conditions is expected to grow exponentially – a hallmark of Lyapunov instability, which is often taken as a defining feature of classical chaos. Quantum mechanically, this exponential growth is expected to be captured by OTOCs in systems that have a well-defined semiclassical limit or large-$N$ limit. $D_{k,\ell}(\tau\to\infty)\ll 1$ corresponds to situations where the phases in replica $a$ and $b$ remain strongly correlated, suggestive of regular dynamics. In contrast, in chaotic regimes where the trajectories decorrelate rapidly, $\varphi^{(a)}$ and $\varphi^{(b)}$ can be seen as uniformly distributed random phases, resulting in a saturation value at its ergodic bound $D_{k,\ell}(\tau\to\infty)\approx 1$. We refer the Reader to the Methods section for additional details on the OTOC’s dynamics. Here, we use the saturation value of the steady-state phase OTOC, $D_{1,\ell}(\tau\to\infty)$, as a proxy to map the chaotic versus regular character of the nonequilibrium steady-state dynamics as the drive strength $F$ and the chain length $L$ are varied. We notice that, while $n_{\ell}$ and $\delta n_{\ell}$ (or $g^{(2)}_{\ell}$) can be monitored in the laboratory, accessing the behavior of $D_{k\ell}$ may be more difficult, as it requires an accurate cloning of the time evolution.
[…]
@@
Focusing on the last site, $\ell=L$, the results are collected in Fig. 2. Qualitatively similar results are obtained throughout the chain, and we refer the reader to the Supplementary Information for results at $\ell=1$ and $\ell=L/2$. We identify three distinct regimes:
@@
    Regular quasilinear regime; at weak drive, only a few photons populate the last site. The photon statistics are Poissonian ($\delta n_{L}\approx 0$) and the small saturation value of the phase OTOC, $D_{1,L}(\tau\to\infty)\approx 10^{-2}-10^{-1}$, indicates regular dynamics. In this regime, single-particle excitations are dilute, rendering nonlinearities negligible and preventing the onset of chaos.
@@
    Chaotic regime; at stronger drives, the photon number markedly increases with respect to the quasilinear regime and nonlinearities are now relevant. The large value of $\delta n_{L}$ indicates strong super-Poissonian fluctuations, compatible with a thermal description of light. Concurrently, the phase OTOC saturates to $D_{1,L}(\tau\to\infty)\approx 1$, signaling significant dephasing between the fields in the two replicas. These observations, along with the exponential growth of $D_{1,\,\ell}(\tau)$ shown in the Methods section, confirm the chaotic nature of the quantum dynamics. This regime will be discussed in detail in Sect. II.3.
@@
    Resonant nonlinear wave (RNW) regime; at even stronger drives, the dynamics reach an incompressible regime where the large photon number marginally increases with $F$ [80]. The small saturation value $D_{1,\,L}(\tau\to\infty)\approx 10^{-2}-10^{-1}$ indicates persistent phase correlations between the two replicas. These observations, along with the sub-exponential growth of $D_{1,\,\ell}(\tau)$ shown in the Methods section, are indicative of regular dynamics. This regime exhibits distinct quantum signatures, most notably the emergence of sub-Poissonian photon statistics, as indicated by $\delta n_{L}<0$. These features will be discussed in detail in Sect. II.4.
Notably, the crossover from the regular quasilinear regime (I) to the chaotic regime (II) is smooth in both $n_{L}$ and $\delta n_{L}$ while the transition between the chaotic regime (II) and the RNW regime (III) is characterized by abrupt variations [see Figs. 2 (d) and (e)]. In particular, the drive strength $F$ at which the photon number leaps coincides with that of the maximum of $\delta n_{L}$.
II.3 Chaotic regime
II.3.1 Two-stage relaxation in space
In this section, we focus on the chaotic regime (II). We decompose the complex field in terms of amplitude and phase degrees of freedom, $\alpha_{\ell}=|\alpha_{\ell}|\,{\rm e}^{{\rm i}\varphi_{\ell}}$, and separately analyze their spatial relaxation along the chain, from $\ell=1$ to $\ell=L$. Figures 3 (a) and (b) shows the steady-state profile of photon density $n_{\ell}=\langle|\alpha_{\ell}|^{2}\rangle-1/2$ and the circular phase variance $\Delta\varphi_{\ell}:=1-|\langle{\rm e}^{{\rm i}\varphi_{\ell}}\rangle|$ across the chain. The latter quantifies the spread of the phase distribution: $\Delta\varphi_{\ell}\to 0$ if the phase is well-defined, whereas $\Delta\varphi_{\ell}\to 1$ for a uniformly distributed phase. The field amplitude decays slowly along the chain over a characteristic length scale $\xi_{n}$, which increases monotonically with both the chain length $L$ and the drive strength $F$. We distinguish two regimes: for weak drives, $F\lesssim 6.5$, the photon density relaxes just beyond the driven site; for intermediate drives, $F\gtrsim 6.5$, the density first saturates to a large value over an extended region, with its decay occurring only near the end of the chain. These large spatial scales can be explained by the local conservation of the photon number within the bulk of the chain, which hinders the rapid relaxation of the amplitude degree of freedom. Instead, this relaxation occurs through much slower hydrodynamic processes involving, notably, the driven-dissipative conditions at the two boundaries of the chain. In contrast, the phase degrees of freedom relax over a microscopic length scale $\xi_{\varphi}$, typically spanning only a few lattice sites. The saturation of the phase variance, $\Delta\varphi_{\ell}\to 1$, signals the restoration of the $\mathbb{U}(1)$ symmetry of the underlying bulk Hamiltonian. This mechanism is illustrated in Figs. 3 (d-f), where the local Wigner functions $W_{\ell}(\alpha,\alpha^{*})$ rapidly lose the anisotropy imprinted by the coherent boundary drive and become invariant under rotations in the complex plane throughout the rest of the chain. The overall change of the topology of $W_{\ell}(\alpha,\alpha^{*})$, from ring shape to bell shape, will be discussed below.
@@ II.3.1 Two-stage relaxation in space
$g^{(1)}_{k,\ell}:=\frac{\langle\hat{a}_{k}^{\dagger}\hat{a}_{\ell}\rangle}{\sqrt{\langle\hat{a}_{k}^{\dagger}\hat{a}_{k}\rangle\langle\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\rangle}}\,.$ (7)
@@
The results are presented in Fig. 3 (c) as a function of $\ell$ for a center site $k=L/2$. The profile of the spatial correlations shows exponential decay, $|g^{(1)}_{k,\ell}|\sim{\rm e}^{-|k-\ell|/\xi_{\varphi}}$, typical of a disordered phase with a correlation length $\xi_{\varphi}$ on the order of a few sites, before saturating to values below $10^{-2}$.
@@
Altogether, these results point to a two-stage thermalization process in space: the phase sector relaxes to a disordered state over a short, microscopic length scale, while the amplitude sector undergoes a slower relaxation that depends on the boundary conditions and unfolds over the entire system. As we shall demonstrate below, this opens up a significant region of the chain, beyond the few sites directly influenced by the boundary drive, but before the complete relaxation to near-vacuum states subject to thermal fluctuations, where an effective $\mathbb{U}(1)$ symmetry emerges. In this intermediate region, the highly populated sites along the chain can be interpreted as being in local equilibrium, characterized by a large and slowly varying chemical potential.
@@
Figure 4: Chaotic regime: hydrodynamic description. Spatial profile of the local (a) effective temperature $T_{\ell}$, (b) effective chemical potential $\mu_{\ell}$, and (c) entropy density $S_{\ell}:=-\textrm{Tr}[\hat{\rho}_{\ell}\log\hat{\rho}_{\ell}]$ across a chain of length $L=400$. The results are obtained by mapping the local steady-state physics at site $\ell$ to the Gibbs ansatz defined in Eq. (8). (d) Spatial profile of the dimensionless ratio $\mu_{\ell}/T_{\ell}$ for the same system. The results are obtained by mapping the local physics at site $\ell$ to the Gibbs ansatz, to the 2-photon impurity ansatz defined in Eqs. (9-10), and to the generalized Scully-Lamb ansatz. Three distinct spatial domains are discussed in the text: 1) nonsymmetric nonthermal domain where all the ansätze fail; 2) prethermal domain where $\mu_{\ell}/T_{\ell}>0$; 3) thermal domain with $\mu_{\ell}/T_{\ell}\leq 0$. (e) Respective sizes of the three domains versus the total chain length $L$. The drive strength is set to $F=7.5$ and the other parameters are set as in Fig. 2.
II.3.2 Hydrodynamic description
@@
We test this hydrodynamic scenario by probing the local symmetries, the local thermodynamics, and the temporal fluctuations across the chain. On general grounds, the local steady-state density matrix $\hat{\rho}_{\ell}:=\lim\limits_{t\to\infty}\textrm{Tr}_{k\neq\ell}\hat{\rho}(t)$ can be cast as the solution of an effective driven-dissipative impurity model constructed by singling out one site of the chain, and tracing out its neighbors such as to treat them as incoherent sources and drains. In a generic NESS, however, the rigorous derivation of an effective driven-dissipative impurity model is a daunting task, and we simply rely on symmetries and heuristics to identify minimal impurity models that successfully locally capture both the statics as well as temporal fluctuations.
Let us first address the statics. Guided by hydrodynamic principles, we assume that, away from the boundary drive, the local steady-state density matrix $\hat{\rho}_{\ell}$ is close to an equilibrium $\mathbb{U}(1)$-symmetric Gibbs state of the form (we set $k_{\rm B}=1$)
$\hat{\rho}^{\rm eq}_{\ell}=\exp[-(\hat{h}-\mu_{\ell}\,\hat{a}^{\dagger}\hat{a})/T_{\ell}]/\mathcal{Z}_{\ell}\,,$ (8)
where $\mathcal{Z}_{\ell}$ is such that $\mathrm{Tr}\left(\hat{\rho}^{\rm eq}_{\ell}\right)=1$, and the local Hamiltonian
$\hat{h}=\frac{1}{2}U\,(\hat{a}^{\dagger}\hat{a})^{2}\,$ (9)
corresponds to a single site of the chain Hamiltonian in Eq. (2). $T_{\ell}$ and $\mu_{\ell}$ are, respectively, the effective temperature and chemical potential at site $\ell$, to be determined. Note that the quadratic onsite contributions to $\hat{H}$, namely $-\Delta-U/2$, have been absorbed into a redefinition of $\mu_{\ell}$.
We assess the validity of this effective description and the underlying restoration of the $\mathbb{U}(1)$ symmetry by fitting the local steady-state Wigner function $W_{\ell}(\alpha,\alpha^{*})$ along the chain to those predicted by the above Gibbs ansatz. The two fitting parameters are $T_{\ell}$ and $\mu_{\ell}$. We repeat this procedure at all sites from $\ell=1$ to $\ell=L$. Excellent matches are obtained at all sites except near the driven boundary, thereby confirming the rapid restoration of the $\mathbb{U}(1)$ symmetry away from the drive and validating the use of the Gibbs state as a local thermometer. In regimes where the photon density is very low, the Kerr nonlinearity is effectively inactive, and it becomes numerically challenging to extract $T_{\ell}$ and $\mu_{\ell}$ independently. However, the ratio $\mu_{\ell}/T_{\ell}$, which governs the shape of the Wigner function in the dilute limit, can still be determined with high accuracy.
Let us now turn to the description of the dynamical fluctuations in the steady state. We start from the local $\mathbb{U}(1)$-symmetric Hamiltonian $\hat{h}$ introduced in Eq. (9) and promote it to Lindblad dynamics by supplementing it with the first non-trivial jump operators allowed under weak $\mathbb{U}(1)$ symmetry [81],
$\hat{L}_{\uparrow}\!=\!\sqrt{\gamma^{\uparrow}_{\ell}}\hat{a}^{\dagger}\!,\hat{L}_{\downarrow}\!=\!\sqrt{\gamma^{\downarrow}_{\ell}}\hat{a},\hat{L}_{\phi}\!=\!\sqrt{\gamma^{\phi}_{\ell}}\hat{a}^{\dagger}\hat{a},\hat{L}_{\rm s}\!=\!\sqrt{\gamma^{\rm s}_{\ell}}\hat{a}^{2}.$ (10)
@@
The non-negative parameters $\gamma^{\uparrow}_{\ell}$, $\gamma^{\downarrow}_{\ell}$, $\gamma^{\phi}_{\ell}$, and $\gamma^{\rm s}_{\ell}$ are effective rates of incoherent pumping, decay, dephasing, and 2-photon decay at site $\ell$. The inclusion of the latter is essential to ensure the saturation of the photon number when $\gamma^{\uparrow}_{\ell}>\gamma^{\downarrow}_{\ell}$. Equations (9) and (10) define a driven-dissipative impurity ansatz that supports a NESS. The parameters $\gamma^{\uparrow}_{\ell}$, $\gamma^{\downarrow}_{\ell}$ and $\gamma^{\rm s}_{\ell}$ are determined by fitting the local Wigner functions $W_{\ell}(\alpha,\alpha^{*})$ to those predicted by the impurity ansatz. The latter are independent of the dephasing rate $\gamma_{\ell}^{\phi}$ that can be extracted by fitting two-time correlation functions. The overall fitting procedure is detailed in the Supplementary Information. Excellent matches are obtained everywhere away from the boundary drive.
Although this impurity ansatz does not rely on assuming local equilibria, we have verified that both the previous Gibbs ansatz and this impurity ansatz consistently capture and reproduce the same static properties when related through the detailed-balance condition
$\mu_{\ell}/T_{\ell}=\log(\gamma^{\uparrow}_{\ell}\big{/}\gamma^{\downarrow}_{\ell})\,.$ (11)